OpenStudy (anonymous):

please help!!!!! the length of a rectangle is 5 yd less than twice the width, and the area of the rectangle is 33 yd^2. find the dimensions of a rectangle.

OpenStudy (jennyrlz):

heyo :)

OpenStudy (jennyrlz):

are yous still here?

OpenStudy (jennyrlz):

hmm, would prefer to work it out with you but i guess i can just show my work if you are not here c:

OpenStudy (jennyrlz):

ok so we know the equation to find the are of a rectangle is: Height x Width

OpenStudy (jennyrlz):

now there are two ways to go about doing this, the "safe" way to set it up into a n equation. While this method takes longer you will have less of a chance of making mistakes!

OpenStudy (jennyrlz):

so we can set it up like x-5=2x

OpenStudy (jennyrlz):

2x-5=x*

OpenStudy (jennyrlz):

hmm i ended up mixing my self up xD

OpenStudy (jennyrlz):

let me show you the carless way i did it in my head ^.^;

OpenStudy (jennyrlz):

33

OpenStudy (jennyrlz):

then by simple multiplication or division, i instintly arived at 11 x 3

OpenStudy (jennyrlz):

3 x 2 = 6 6 + 5 =11 or 11 - 5 = 6

OpenStudy (jennyrlz):

hmm i cant seem to find out how to put it into an equation 0.o

OpenStudy (jennyrlz):

sorry xD

OpenStudy (jennyrlz):

but i am sure of the answer though :)